Trade Winds Enterprises Ltd (TW) has £20,000 that it can invest in any or all of the two capital investment projects, which have cash flows as shown b Comparison of Project Cash Flows Year of Cash Flow Year 0 Year l Type of Cash Flow Year 3 Poject pe of Year 2 (E10 000) Investment Revenue Operating expenses Investment Revenue Operating expenses A. £10000 £1000 £30 000 4889 15 555 (10 000) 30000 10000 2 222 5 555 2 222 5 55S B. 15 555 The investment will be depreciated to zero on a straight-line basis for tax purposes. TWs marginal corporate tax rate on taxable income is 40%. None of the projects will have any salvage value at the end of their respective lives. For purposes of analysis, it should be assumed that all cash flows occur at the end of the year in question. Rank TWs two projects according to the following three commonly used capital budgeting criteria: Payback period. Accounting return on investment (ROI). Net present value (NPV), assuming alternately a 10% discount rate (Present Value and Annuity tables are provided) . If the projects are independent of each other, which should be accepted? If they are mutually exclusive (i.e., one and only one can be accepted), which one is best?
Answer #1

ANSWER:

Payback Periods:

PROJECT A:

(Table drawn on paper)

PnojcatA ye Cash ca) b e pevesesPayback Period = Initial Investment / cashflow to cover initial investment.

= 1 year + (5555/6111)

= 1+ 0.909 year

= 2 years approx.

PROJECT B:

(Table drawn on paper)

Proje tB Ca to COL ) Cann 3°.cso | ls,SS. S | ( 나,4 4 Yİ 4445Payback Period = Initial Investment / cash flow to cover initial investment.

= 10000 / 14445

= 0.69 year

= 7 months approx.

Conclusion: Project B has short payback period in comparison to Project A. Therefore, Project B is profitable

ACCOUNTING RETURN ON INVESTMENT:

Accounting return on investment = Average net profit / Average investment

Average net profit = Estimated net average annual income after depreciation and tax

PROJECT A:

(Table on Paper)

Re Ventue Cb) C 3 333. 337 68 6 66 total-7ooo

Depreciation (SLM) = (Original value - Scarp value ) / Estimated life of the asset

= (10000 - 0) / 3

= 3333.33

Accounting return on investment = Average net profit / Average investment

= 3000 / 10000

= 0.3

Average net profit = 9000 / 3 = 3000

PROJECT B:

(table on paper)

Project Deprecisctinevenue 1444r |召333.33 | 오1111.67 | tu44.64S | CCC 구 3333 -33 33 SS.33 2다귀8

Depreciation (SLM) = (Original value - Scarp value ) / Estimated life of the asset

= (10000 - 0) / 3

= 3333.33

Accounting return on investment = Average net profit / Average investment

= 2260/ 10000

= 0.226

. Average net profit = 6778.67/ 3 = 2260